# https://jeonghyeokdo.github.io/SkateFormer_site/

### v1

@@ 4.2 Experiment Details
From various experiments, we found the empirical values as follows. For our experiments with the NTU RGB+D and NTU RGB+D 120 datasets, we used the following configuration: $V=48$ (excluding the center of the body, resulting in 24 joints for each of the two individuals), $T=64$, $L=4$, $K=12$, $M=8$, $N=8$, and $C=96$. For the experiments with the NW-UCLA dataset, $V=20$, $T=64$, $L=4$, $K=5$, $M=8$, $N=8$, and $C=96$. Furthermore, we employed a total of $R=8$ SkateFormer Blocks, applying temporal downsampling after every 2 blocks. The kernel size for the $\mathsf{T{\text{-}}Conv}$ layers was set to $k=7$. For the performance evaluation, we compare our SkateFormer with other methods in terms of the top-1 action recognition accuracy for the test data sets.
@@ 4.4 Ablation Studies
Table 4: Ablation study on the influence of skeletal and temporal relation types (SkateTypes) of Skate-MSA in our SkateFormer.
  Relation Types NTU RGB+D (%) Params. (M) FLOPs (G) Time (ms)
@@
Baseline (only $\mathbf{V}_{h}$ in Eq. 8) 90.7 95.7 2.03 3.591 10.68
Baseline + $\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ (skeletal) 91.8 ($\uparrow$1.1) 96.4 ($\uparrow$0.7) 2.03 3.594 11.10
Baseline + $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$ (temporal) 91.9 ($\uparrow$1.2) 96.6 ($\uparrow$0.9) 2.03 3.593 11.09
Baseline + $\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ + $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$ 92.6 ($\uparrow$1.9) 97.0 ($\uparrow$1.3) 2.03 3.620 11.46
@@ 0.B.1 Recognition Accuracy based on Action Labels
Table 8 presents the top-1 accuracy results for single-joint modality (J) action recognition in the NTU RGB+D X-Sub60 evaluation, categorized by action labels. The baseline model represents a model without the application of partition-specific attention based on four skeletal-temporal relation types from Skate-MSA (only $\mathbf{V}_{h}$ in Eq. 8). The model denoted as (+ $\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$) applies partition-specific attention based on two skeletal relation types, while (+ $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$) utilizes partition-specific attention based on two temporal relation types. The model labeled (+$\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ + $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$) represents the full SkateFormer model, which incorporates partition-specific attention based on four Skate-Types.
@@ 0.F.1 Hyperparameters
Here is a more detailed description of the hyperparameters we used. We employed the AdamW optimizer [32] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay value of 0.1. For the cosine-annealing scheduler [31], we set warmup learning rate to $10^{-7}$, minimum learning rate to $10^{-5}$, base learning rate to $10^{-3}$, warm-up epochs to 25, and utilized linear warm-up strategy. To prevent overfitting in the network, we set the attention map dropout ratio to 0.5 and the drop path ratio to 0.2. We did not use dropout for classification head and linear layers. The number of heads $H$ was consistent across all SkateFormer Blocks, set at 32. The expansion ratio of FFN was set to 1 for the NW-UCLA [56] dataset and 4 for the NTU RGB+D [41] and NTU RGB+D 120 [27] datasets. We employed the GELU activation function in our model.

### v3

# https://kaist-viclab.github.io/SkateFormer_site/

@@ 4.2 Experiment Details
From various experiments, we found the empirical values as follows. For our experiments with the NTU RGB+D and NTU RGB+D 120 datasets, we used the following configuration: $V=48$ (excluding the center of the body, resulting in 24 joints for each of the two individuals), $T=64$, $L=4$, $K=12$, $M=8$, $N=8$, and $C=96$. For the experiments with the NW-UCLA dataset, $V=20$, $T=64$, $L=4$, $K=5$, $M=8$, $N=8$, and $C=96$.
@@ 4.4 Ablation Studies
Table 4: Ablation study on the influence of skeletal and temporal relation types (SkateTypes) of Skate-MSA in our SkateFormer. (*: The batch size is set to 8 due to limited GPU memory.)
  Attention Types NTU RGB+D (%) Params. (M) FLOPs (G) Time (ms) Memory (MB)
@@
Baseline (no attention) 90.7 95.7 2.03 3.59 10.68 80.6
+ Naive self-attention* 90.6 ($\downarrow$0.1) 95.0 ($\downarrow$0.7) 2.03 4.39 17.55 1612.1
+ $\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ 91.8 ($\uparrow$1.1) 96.4 ($\uparrow$0.7) 2.03 3.59 11.10 86.9
+ $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$ 91.9 ($\uparrow$1.2) 96.6 ($\uparrow$0.9) 2.03 3.59 11.09 85.5
+ $\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ + $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$ 92.6 ($\uparrow$1.9) 97.0 ($\uparrow$1.3) 2.03 3.62 11.46 137.6
@@ 5 Conclusions
Acknowledgements. This work was supported by IITP grant funded by the Korea government(MSIT) (No.RS2022-00144444, Deep Learning Based Visual Representational Learning and Rendering of Static and Dynamic Scenes).
Supplementary Material
  Supple. Sec. Descriptions for analysis and discussion
Sec. 0.A High-level novelty compared to conventional methods
Sec. 0.B.1 Statistical analysis of Skate-MSA for different classes
(i) Accuracy comparisons of SkateFormer for different classes (Table. 8) (ii) Statistical analysis of action classes where Skate-MSA enhances performance (iii) Discussion of failure action classes
Sec. 0.B.2 Generalization issues to diverse modality inputs
Sec. 0.B.3 Visualizing the activation levels of Skate-MSA
(i) Visualization of activation for different classes (Fig. 6) (ii) Discussion of adaptive nature of Skate-MSA concerning action classes
Sec. 0.E.1 Theoretical calculation of the computational complexity of Skate-MSA
Sec. 0.E.2 Additional experiments on the computational complexity
Due to the space limit of the main paper, we had a difficulty to contain all the detailed descriptions for the analysis and discussion on the various experimental results in the same granularity levels of details. Instead, high level descriptions are included in the main paper while detailed descriptions are in the Supple. as shown in the above table.
@@ 0.B.1 Recognition Accuracy based on Action Labels
Table 8 presents the top-1 accuracy results for single-joint modality (J) action recognition in the NTU RGB+D X-Sub60 evaluation, categorized by action labels. The baseline model represents a model without the application of partition-specific attention based on four skeletal-temporal relation types from Skate-MSA (no attention, only $\mathbf{V}_{h}$ in Eq. 8). The model denoted as (+ $\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$) applies partition-specific attention based on two skeletal relation types, while (+ $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$) utilizes partition-specific attention based on two temporal relation types. The model labeled (+$\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ + $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$) represents the full SkateFormer model, which incorporates partition-specific attention based on four Skate-Types.
@@ 0.E.1 Computational Complexity of Skate-MSA
0.E.2 Additional Experiments on the Computational Complexity
Our SkateFormer comprises 8 SkateFormer blocks and 3 temporal downsampling layers. The 48$\times$ reduction in FLOPs was theoretically calculated from the first two SkateFormer blocks before downsampling, where the reduction effect is most pronounced. To clarify this, we additionally performed experiments in Table 12 in terms of measured FLOPs and memory usages in perspective of the attention layers only. As shown, the Skate-MSA is 38.87$\times$ less complex in terms of FLOPs for the first two attention layers.
Table 12: FLOPs and memory usage for attention layers.
  Attention Types in attention layer Total 8 attention layers First two attention layers
FLOPs (M) Memory (MB) FLOPs (M) Memory (MB)
$\mathbf{v}_{k}^{\mathsf{njp}}$ + $\mathbf{v}_{l}^{\mathsf{djp}}$ + $\mathbf{t}_{m}^{\mathsf{local}}$ + $\mathbf{t}_{n}^{\mathsf{global}}$ 29.12 56.3 15.53 27.0
Naive self-attention 801.79 ($\times$27.53) 1530.0 ($\times$27.18) 603.65 ($\times$38.87) 1152.0 ($\times$42.67)
@@ 0.F.1 Hyperparameters
Here is a more detailed description of the hyperparameters we used. We employed the AdamW optimizer [32] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay value of 0.1. For the cosine-annealing scheduler [31], we set warmup learning rate to $10^{-7}$, minimum learning rate to $10^{-5}$, base learning rate to $10^{-3}$, warm-up epochs to 25, and utilized linear warm-up strategy. To prevent overfitting in the network, we set the attention map dropout ratio to 0.5 and the drop path ratio to 0.2. We did not use dropout for classification head and linear layers. The number of heads $H$ was consistent across all SkateFormer Blocks, set at 32. The expansion ratio of FFN was set to 1 for the NW-UCLA [56] dataset and 4 for the NTU RGB+D [41] and NTU RGB+D 120 [27] datasets. We employed the GELU activation function in our model. Furthermore, we employed a total of $R=8$ SkateFormer Blocks, applying temporal downsampling after every 2 blocks. The kernel size for the $\mathsf{T{\text{-}}Conv}$ layers was set to $k=7$. For the performance evaluation, we compare our SkateFormer with other methods in terms of the top-1 action recognition accuracy for the test data sets.